3) The following relations are given for an economy:
Autonomous Consumption 460
Marginal Propensity to Save 0.3
Autonomous Investment 18
Induced Investment Coefficient 0.1
Government Expenditure (Exogenous) 162
Transfer Income (Exogenous) 48
Exports (Exogenous) 25
Marginal Propensity to Import 0.05
a) Calculate the equilibrium level of Income, consumption, investment and import in the
above economy.
1
Expert's answer
2020-04-27T07:54:38-0400

Y = C+I+G

Marginal Propensity to Import = import/Y

Import = 0.05 Y

G=162

C=460+0.7Y

investment = 18 +0.1Y

Y= 460+0.7 Y+18+0.1Y+162

Y= 0.8Y+640

0.2Y=640

Y=3200

Import = 0.05*3200

Import= 160